OpenStudy (lin.ivory):

Probability question! A box contains one yellow, two red, and three green balls. Two balls are randomly chosen without replacement. A: {One of the balls is yellow} B: {At least one ball is red} C: {Both balls are green} D: {Both balls are of the same color} Find: P(A|B-compliment) P(A-overline|B)

OpenStudy (kropot72):

\[P(1\ yellow)=\] \[P(1\ yellow)=\frac{\left(\begin{matrix}1 \\ 1\end{matrix}\right)\left(\begin{matrix}5 \\ 1\end{matrix}\right)}{\left(\begin{matrix}6 \\ 2\end{matrix}\right)}=\frac{5}{15}\]

OpenStudy (kropot72):

\[P(1\ red)=\frac{2C1\times 4C1}{6C2}=you\ can\ calculate?\] \[P(2\ red)=\frac{2C2\times 4C0}{6C2}=you\ can\ calculate?\] \[P(at\ least\ 1\ red)=P(1\ red)+P(2\ red)\]

OpenStudy (lin.ivory):

but 5/15 is not the answer :(

OpenStudy (kropot72):

But the answer to A must be simplified as follows giving two possible forms of the answer: \[\frac{3}{15}=\frac{1}{3}\] Or alternatively \[\frac{3}{15}=0.3333\]

OpenStudy (lin.ivory):

ohh!! that makes sense now! thanks!

OpenStudy (kropot72):

You're welcome :)